# Linear and nonlinear instability of vortex columns

### v2

@@ 3.1. The Biot-Savart law
The Biot-Savart law $u={\rm BS}[\omega]$ requires special attention on the slab domain $\mathbb{R}^{2}\times\mathbb{T}$ since, when $\omega$ is independent of $z$, the problem is two dimensional. [^12] The problem is more precisely two-and-a-half dimensional. The two-dimensional Biot-Savart law is well defined on $m$-fold rotationally symmetric vorticities in $L^{2}$, $m\geq 2$, as exploited in [53, 4]. The two-and-a-half dimensional case also contains vertical shears like $u=f(x^{\prime})e_{z}$, $x^{\prime}\in\mathbb{R}^{3}$, with vorticity $\omega=(\nabla^{\perp}_{x^{\prime}}f,0)$ and $u=\Delta^{-1}_{x^{\prime}}\nabla^{\perp}_{x^{\prime}}\cdot\omega$. The two-dimensional Biot-Savart law maps $L^{2}\to\dot{H}^{1}$ and a priori is only well defined up to constants unless additional decay is assumed; this is precisely the reason for our assumption $\omega\in L^{Q}$, $Q\in(1,2)$.
@@
$\|{\rm BS}[P_{\neq 0}\omega]\|_{W^{1,q}(\mathbb{R}^{2}\times\mathbb{T})}\lesssim\|\omega\|_{L^{q}}\,,\quad q\in(1,2]\,,$ (3.1)
@@
where $p_{2}^{*}:=2p/(2-p)$ denotes the $2$D Sobolev exponent.
@@
$-\Delta\psi=P_{\neq 0}\omega\quad\text{ in }\mathbb{R}^{2}\times\mathbb{T}\qquad\text{ with }\qquad\int_{\mathbb{T}}\psi(x^{\prime},y_{3})\,\mathrm{d}y_{3}=0\quad\text{ for all }x\,.$ (3.3)
@@
Indeed, given (3.4), we can apply the local regularity theory for $-\Delta\psi=P_{\neq 0}\omega$ to demonstrate
@@
In order to prove (3.4), we multiply (3.3) by $|\psi|^{q-2}\psi/(q-1)$ and integrate by parts to obtain
@@
$\int_{B_{1}\times\mathbb{T}}|\nabla|\widetilde{\psi}_{k}|^{q/2}|^{2}+\int_{B_{1}\times\mathbb{T}}|\widetilde{\omega}_{k}|^{q}\to 0\,.$ (3.12)
By the local regularity theory, we moreover have
@@
By the compactness granted in (3.9), we have
$|\widetilde{\psi}_{k}|^{q/2}\to|\widetilde{\psi}_{\infty}|^{q/2}\text{ in }L^{2}\,,\quad\nabla|\widetilde{\psi}_{\infty}|^{q/2}=0\,,\quad\||\widetilde{\psi}_{\infty}|^{q/2}\|_{L^{2}}=1\,,$ (3.16)
which implies that $|\widetilde{\psi}_{\infty}|^{q/2}$, and therefore $|\widetilde{\psi}_{\infty}|$ is a non-zero constant. Since $\widetilde{\psi}_{\infty}$ solves $\Delta\widetilde{\psi}_{\infty}=0$, it is a smooth function; since it is moreover scalar-valued, $\widetilde{\psi}_{\infty}$ is therefore a non-zero constant. Finally, since $\widetilde{\psi}_{\infty}$ has zero mean in $x_{3}$, it must be zero. This contradicts that $\|\widetilde{\psi}_{\infty}\|_{L^{q}}=1$ and completes the proof of (3.8).
@@
The horizontal velocity $(u_{1},u_{2})$ is determined by the 2D Biot-Savart law, and $u_{3}$ is obtained from $(\omega_{1},\omega_{2})$ by $-1$-order Fourier multipliers with homogeneous symbol which is smooth away from the origin. The 2D Calderón-Zygmund inequality and the Sobolev inequality therefore yield the desired estimate (3.2) when $\omega\in L^{p}$ and $21<p<$ 2. ∎
@@ 3.2. Properties of ${\mathcal{L}}$
We study properties of the operator ${\mathcal{L}}$, defined as a linearization of the $3$D Euler equations (1.1) around ${\overline{u}}$. Recall from (1.37) that ${\mathcal{L}}:=D({\mathcal{L}})\subset X\to X$ is defined by
@@
$(\lambda-{\mathcal{L}})\omega=f\,,$ (3.19)
@@
$(|\mathrm{Re}\lambda|-\|{\widetilde{L}}\|_{L^{Q}_{\sigma}\to L^{Q}_{\sigma}})\|\omega\|_{L^{Q}_{\sigma}}\leq\|f\|_{L^{Q}_{\sigma}}\,,$ (3.20)
obtained when we multiply (3.19) by $|\omega|^{Q-2}\omega$ and integrate by parts. Existence can be proved, for example, by considering vanishing viscosity approximations. We emphasize that (3.19) involved ${\mathcal{L}}$ (rather than ${\widetilde{L}}$), and the above argument makes use of the particular structure of ${\overline{u}}\cdot\nabla(\cdot)$.
@@
Thus $\omega+{\overline{\omega}}$ satisfies the $3$D Euler equations if
@@
We recover ${\mathcal{L}}$ by considering only the last two terms in the first line of (3.22). Namely, writing [^14] By this, we mean that $L^{Q}_{\sigma}$ is the $L^{Q}$-closure of finite linear combinations of elements of $\cup_{n,\alpha}X_{n,\alpha}e^{i(\alpha z+n\theta)}$.
@@
$-A_{n,\alpha}\begin{pmatrix}w_{r}\\ w_{\theta}\\ w_{z}\end{pmatrix}:=\begin{pmatrix}in\frac{V}{r}+i\alpha W&0&0\\ -V^{\prime}+\frac{V}{r}&in\frac{V}{r}+i\alpha W&0\\ -W^{\prime}&0&in\frac{V}{r}+i\alpha W\end{pmatrix}\begin{pmatrix}w_{r}\\ w_{\theta}\\ w_{z}\end{pmatrix}$ (3.23)
$-B_{n,\alpha}\begin{pmatrix}w_{r}\\ w_{\theta}\\ w_{z}\end{pmatrix}:=\begin{pmatrix}-in\frac{W^{\prime}}{r}-i\alpha\frac{(rV)^{\prime}}{r}&0&0\\ W^{\prime\prime}-\frac{W^{\prime}}{r}&-in\frac{W^{\prime}}{r}-i\alpha\frac{(rV)^{\prime}}{r}&0\\ \left(\frac{(rV)^{\prime}}{r}\right)^{\prime}&0&-in\frac{W^{\prime}}{r}-i\alpha\frac{(rV)^{\prime}}{r}\end{pmatrix}\begin{pmatrix}u_{r}\\ u_{\theta}\\ u_{z}\end{pmatrix}\,.$ (3.24)
@@
For (i), we observe that $A_{n,\alpha}$ is multiplication operator by a bounded matrix function. Clearly, $\lambda-A_{n,\alpha}$ is invertible when $\lambda$ is away from the range of $-in\Omega-i\alpha W$, since the determinant of $\lambda-A_{n,\alpha}$ will be bounded away from zero. When $\lambda=-in\Omega(r_{0})-i\alpha W(r_{0})$ is in the range, we simply approximate $\delta_{r_{0}}(r)e_{z}$ by an $L^{Q}$ function. Finally, recall that the spectrum is closed.
@@
$\begin{split}\omega\cdot\nabla{\overline{u}}&=\mathrm{e}^{in\theta+i\alpha z}(w_{r}e_{r}+w_{\theta}e_{\theta}+w_{z}e_{z})\cdot\nabla(Ve_{\theta}+We_{z})\\ &=\mathrm{e}^{in\theta+i\alpha z}(w_{r}\partial_{r}+w_{\theta}\frac{\partial_{\theta}}{r}(Ve_{\theta}+We_{z})\\ &=\mathrm{e}^{in\theta+i\alpha z}(w_{r}(V^{\prime}e_{\theta}+W^{\prime}e_{z})-w_{\theta}\frac{V}{r}e_{r})\\ &=\frac{V^{\prime}}{r^{2}}(\omega\cdot x)x^{\perp}+\frac{W^{\prime}}{r}(\omega\cdot x)e_{z}-\frac{V}{r^{3}}(\omega\cdot x^{\perp})x,\end{split}$
@@
$\begin{split}\lambda\omega+\overline{u}\cdot\nabla\omega-\omega\cdot\nabla\overline{u}&=\lambda\omega+\left(-in\frac{V}{r}+i\alpha W\right)\omega+\frac{V}{r^{3}}(\omega\cdot x)x^{\perp}-\left(\frac{V^{\prime}}{r^{2}}(\omega\cdot x)x^{\perp}+\frac{W^{\prime}}{r}(\omega\cdot x)e_{z}\right)\\ &=\lambda\omega+\left(-in\frac{V}{r}+i\alpha W\right)\omega-\frac{\Omega^{\prime}}{r}(\omega\cdot x)x^{\perp}-\frac{W^{\prime}}{r}(\omega\cdot x)e_{z}\\ &=(\lambda-M_{n,\alpha})\omega\,,\end{split}$
@@
$-M_{n,\alpha}:=\begin{pmatrix}in\frac{V}{r}+i\alpha W+\frac{\Omega^{\prime}}{r}xy&\frac{\Omega^{\prime}}{r}y^{2}&0\\ -\frac{\Omega^{\prime}}{r}x^{2}&in\frac{V}{r}+i\alpha W-\frac{\Omega^{\prime}}{r}xy&0\\ -\frac{W^{\prime}}{r}x&-\frac{W^{\prime}}{r}y&in\frac{V}{r}+i\alpha W\end{pmatrix}.$
@@
It will be good to verify that solutions of (1.17) yield solutions $u_{r}\in W^{1,Q}$ to (1.9). Due to the $n^{2}r^{-2}$ singularity of the potential $k$ at $r=0$ and asymptotic size $O(n^{2})$ of $k$ as $r\to+\infty$, we have, starting from the information in Theorem 1.1, that $|u_{r}(r)|\lesssim r^{cn}$ as $r\to 0^{+}$ and $|u_{r}(r)|\lesssim\mathrm{e}^{-cnr}$ as $r\to\infty$ for some constant $c>0$. In particular, $\frac{u_{r}}{r},ru_{r}^{\prime},r^{2}u_{r}^{\prime\prime}\in L^{2}(\mathbb{R}^{+})$, which can be obtained by a simple weighted energy estimate for the equation (1.9). This and (1.13) show that the eigenfunction $u$ of the form (1.8) belongs to $W^{1,Q}(\mathbb{R}^{2}\times\mathbb{T})$, which gives the following.
Corollary 3.4 (Linear instability in $W^{1,Q}$).
Let $M\geq 1$. For every $m\in\{1,\ldots,M\}$ and sufficiently large $n$ there exists $\lambda_{m}\in\mathbb{C}$ with $\mathrm{Re}\,\lambda_{m}>0$ (i.e. $\lambda_{m}=-i\omega_{m}$) and eigenfunction
@@ 3.3. Semigroup bounds for $L_{n,\alpha}$
$\mathrm{e}^{tA_{n,\alpha}}=\mathrm{e}^{-it(n\Omega+\alpha W)}\left(I+t\begin{bmatrix}0&&\\ V^{\prime}+V/r&0&\\ W^{\prime}&&0\end{bmatrix}\right),$ (3.33)
@@
where $A_{n,\alpha}$, $B_{n,\alpha}$ are defined in (3.23)–(3.24). We estimate $B_{n,\alpha}$, which encodes the terms $-\overline{\omega}\cdot\nabla u+u\cdot\nabla\overline{\omega}$. We have
@@
where $1/2=1/Q^{*}_{2}+1/p$. Second, suppose that $Q\in(1,6/5)$. (When $Q\in[6/5,1]$, then $W^{1,Q}\hookrightarrow L^{2}$, which simplifies this estimate.) Then
$\|{\rm BS}[P_{\neq 0}\omega]\cdot\nabla{\overline{\omega}}\|_{L^{2}}\lesssim\|{\rm BS}[P_{0}\omega]\|_{L^{Q}}^{\theta}\|{\rm BS}[P_{0}\omega]\|_{L^{2}}^{1-\theta}\|\nabla{\overline{\omega}}\|_{L^{\infty}}\,,$ (3.37)
@@
[…]

### v3

@@ 3.1. The Biot-Savart law
The Biot-Savart law $u={\rm BS}[\omega]$ requires special attention on the slab domain $\mathbb{R}^{2}\times\mathbb{T}$ since, when $\omega$ is independent of $z$, the problem is two dimensional. [^12] The problem is more precisely two-and-a-half dimensional. The two-dimensional Biot-Savart law is well defined on $m$-fold rotationally symmetric vorticities in $L^{2}$, $m\geq 2$, as exploited in [55, 4]. The two-and-a-half dimensional case also contains vertical shears like $u=f(x^{\prime})e_{z}$, $x^{\prime}\in\mathbb{R}^{2}$, with vorticity $\omega=(-\nabla^{\perp}_{x^{\prime}}f,0)=:(\omega^{\prime},0)$ and $u=(-\Delta_{x^{\prime}})^{-1}(0,\nabla^{\perp}_{x^{\prime}}\cdot\omega^{\prime})$. The two-dimensional Biot-Savart law maps $L^{2}\to\dot{H}^{1}$ and a priori is only well defined up to constants unless additional decay is assumed; this is precisely the reason for our assumption $\omega\in L^{Q}$, $Q\in(1,2)$.
@@
$\|{\rm BS}[P_{\neq 0}\omega]\|_{W^{1,q}}\lesssim\|\omega\|_{L^{q}}\,,\quad\text{ for }q\in(1,2]\,$, (3.1)
@@
where $p_{2}^{*}:=2p/(2-p)$ denotes the 22D Sobolev exponent.
Recall (1.39) that the subscript $\sigma$” in $L^{p}_{\sigma}$ denotes weakly divergence-free vector fields.
@@
$-\Delta\psi=P_{\neq 0}\omega\quad\text{ in }\mathbb{R}^{2}\times\mathbb{T}\qquad\text{ with }$ (3.3)
$\quad\qquad\int_{\mathbb{T}}\psi(x^{\prime},y_{3})\,\mathrm{d}y_{3}=0\text{ and }|\psi(y^{\prime},x_{3})|\to 0\text{ as }|y^{\prime}|\to\infty,\quad\text{ for all }x\,.$
@@
Indeed, given (3.4), we can apply the local regularity theory for $\Delta\psi=P_{\neq 0}\omega$ to demonstrate
@@
In order to prove (3.4), we first multiply (3.3) by $|\psi|^{q-2}\psi/(q-1)$ and integrate by parts to obtain
@@
$\int_{B_{1}\times\mathbb{T}}|\nabla|\widetilde{\psi}_{k}|^{q/2}|^{2}+\int_{B_{1}\times\mathbb{T}}|\widetilde{\omega}_{k}|^{q}\to 0$ (3.12)
as $k\to\infty$. By the local regularity theory, we moreover have
@@
By the compactness granted by (3.11)–(3.12), we have
$|\widetilde{\psi}_{k}|^{q/2}\to|\widetilde{\psi}_{\infty}|^{q/2}\text{ in }L^{2}(B_{1}\times\mathbb{T})\,,\quad\nabla|\widetilde{\psi}_{\infty}|^{q/2}=0\,,\quad\||\widetilde{\psi}_{\infty}|^{q/2}\|_{L^{2}}=1\,$, (3.16)
[…]
@@
The horizontal velocity $(u_{1},u_{2})$ is determined by the 2D Biot-Savart law, and $u_{3}$ is obtained from $(\omega_{1},\omega_{2})$ by 1-1-order Fourier multipliers with homogeneous symbol which is smooth away from the origin. The 2D Calderón-Zygmund inequality and the Sobolev inequality therefore yield the desired estimate (3.2) when $\omega\in L^{p}$ and $1<p<2$. ∎
3.2. Linear instability in $W^{1,Q}$
Here we verify that, after rescaling, the solution $\phi$ constructed in Theorem 1.1 satisfies $u\in W^{1,Q}(\mathbb{R}^{2}\times\mathbb{T})$ for some $Q\in(1,2)$. Eigenfunctions of the form (1.8) are periodic in $z$ with period $2\pi/\alpha$, and we may assume that $\alpha\in 2\pi{\mathbb{N}}$ after rescaling.
For the $W^{1,Q}$ regularity of $u$, we first recall that the potential $k$ in the Rayleigh equation (1.18) admits a singularity of order $n^{2}r^{-2}$ as $r\to 0$, and that $k(r)=O(n^{2})$ as $r\to+\infty$. Thus, from the information on $\phi$ provided by Theorem 1.1, we see that $|u_{r}(r)|\lesssim r^{cn}$ as $r\to 0^{+}$ and $|u_{r}(r)|\lesssim\mathrm{e}^{-cnr}$ as $r\to\infty$ for some constant $c>0$. (Recall also the rescaling (1.17) that $\phi=r^{3/2}(1+\beta^{2}r^{2})^{-1/2}u_{r}$.) In particular, $\frac{u_{r}}{r},ru_{r}^{\prime},r^{2}u_{r}^{\prime\prime}\in L^{2}(\mathbb{R}^{+})$, which can be obtained by a simple weighted energy estimate for the equation (1.9). This and the system (1.13) determining $u_{\theta}$, $u_{z}$ from $u_{r}$ show that the eigenfunction $u$ of the form (1.8) belongs to $W^{1,Q}(\mathbb{R}^{2}\times\mathbb{T})$ for some $Q\in(1,2)$, which gives the following.
Corollary 3.2 (Linear instability in $W^{1,Q}$).
Let $M\geq 1$. For every $m\in\{1,\ldots,M\}$ and sufficiently large $n$ there exist $\lambda_{m}\in\mathbb{C}$ with $\mathrm{Re}\,\lambda_{m}>0$ (i.e. $\lambda_{m}=-i\omega_{m}$) and eigenfunction
@@ 3.3. Properties of $\mathcal{L}$}
We study properties of the operator $\mathcal{L}$}, defined as a linearization of the 33D Euler equations (1.1) around $\overline{u}$}. Recall from (1.38) that ${\mathcal{L}}:=D({\mathcal{L}})\subset X\to X$ is defined by
@@
$(\lambda-{\mathcal{L}})\omega=f$ (3.19)
@@
$\left(|\mathrm{Re}\lambda|-\|{\widetilde{L}}\|_{L^{Q}_{\sigma}\to L^{Q}_{\sigma}}\right)\|\omega\|_{L^{Q}_{\sigma}}\leq\|f\|_{L^{Q}_{\sigma}}\,$, (3.20)
obtained when we multiply (3.19) by $|\omega|^{Q-2}\omega^{*}$ (complex conjugate) and integrate by parts. We emphasize that here we made use of the cancellation
$\int_{\mathbb{R}^{2}\times\mathbb{T}}(\overline{u}\cdot\nabla\omega)\cdot\omega^{*}|\omega|^{Q-2}=0,$
a consequence of the fact that $\mathrm{div}\,\overline{u}=0$, so that the a priori estimate (3.20) involves $\widetilde{L}$} only, while the resolvent problem (3.19) involves $\mathcal{L}$}. We note that the existence of solutions to (3.19) can be proved, for example, by considering vanishing viscosity approximations.
@@
Thus $\omega+{\overline{\omega}}$ satisfies the 33D Euler equations if
@@
We recover $\mathcal{L}$} by considering only the last two terms in the first line of (3.22). Namely, writing [^14] By this, we mean that $L^{Q}_{\sigma}$ is the $L^{Q}$-closure of the set of finite linear combinations of elements of $\bigcup_{n,\alpha}X_{n,\alpha}\mathrm{e}^{i(\alpha z+n\theta)}$.
@@
$A_{n,\alpha}\begin{pmatrix}w_{r}\\ w_{\theta}\\ w_{z}\end{pmatrix}:=\begin{pmatrix}in\frac{V}{r}+i\alpha W&0&0\\ -V^{\prime}+\frac{V}{r}&in\frac{V}{r}+i\alpha W&0\\ -W^{\prime}&0&in\frac{V}{r}+i\alpha W\end{pmatrix}\begin{pmatrix}w_{r}\\ w_{\theta}\\ w_{z}\end{pmatrix}$ (3.23)
$B_{n,\alpha}\begin{pmatrix}w_{r}\\ w_{\theta}\\ w_{z}\end{pmatrix}:=\begin{pmatrix}-in\frac{W^{\prime}}{r}-i\alpha\frac{(rV)^{\prime}}{r}&0&0\\ W^{\prime\prime}-\frac{W^{\prime}}{r}&-in\frac{W^{\prime}}{r}-i\alpha\frac{(rV)^{\prime}}{r}&0\\ \left(\frac{(rV)^{\prime}}{r}\right)^{\prime}&0&-in\frac{W^{\prime}}{r}-i\alpha\frac{(rV)^{\prime}}{r}\end{pmatrix}\begin{pmatrix}u_{r}\\ u_{\theta}\\ u_{z}\end{pmatrix}\,$. (3.24)
@@
For (i), we observe that $A_{n,\alpha}$ is multiplication operator by a bounded matrix function. Clearly, $\lambda-A_{n,\alpha}$ is invertible when $\lambda$ is away from the range of $in\Omega-i\alpha W$, since the determinant of $\lambda-A_{n,\alpha}$ will be bounded away from zero. When $\lambda=-in\Omega(r_{0})-i\alpha W(r_{0})$ is in the range, we simply approximate $\delta_{r_{0}}(r)e_{z}$ by an $L^{Q}$ function. Finally, we recall that the spectrum is closed.
@@
$\begin{split}\omega\cdot\nabla{\overline{u}}&=\mathrm{e}^{in\theta+i\alpha z}(w_{r}e_{r}+w_{\theta}e_{\theta}+w_{z}e_{z})\cdot\nabla(Ve_{\theta}+We_{z})\\ &=\mathrm{e}^{in\theta+i\alpha z}(w_{r}\partial_{r}+w_{\theta}\frac{\partial_{\theta}}{r}(Ve_{\theta}+We_{z})\\ &=\mathrm{e}^{in\theta+i\alpha z}\left(w_{r}(V^{\prime}e_{\theta}+W^{\prime}e_{z})-w_{\theta}\frac{V}{r}e_{r}\right)\\ &=\frac{V^{\prime}}{r^{2}}(\omega\cdot x)x^{\perp}+\frac{W^{\prime}}{r}(\omega\cdot x)e_{z}-\frac{V}{r^{3}}(\omega\cdot x^{\perp})x,\end{split}$
@@
[…]
@@
$M_{n,\alpha}:=\begin{pmatrix}in\frac{V}{r}+i\alpha W+\frac{\Omega^{\prime}}{r}xy&\frac{\Omega^{\prime}}{r}y^{2}&0\\ -\frac{\Omega^{\prime}}{r}x^{2}&in\frac{V}{r}+i\alpha W-\frac{\Omega^{\prime}}{r}xy&0\\ -\frac{W^{\prime}}{r}x&-\frac{W^{\prime}}{r}y&in\frac{V}{r}+i\alpha W\end{pmatrix}$.
@@ 3.4. Semigroup bounds for $L_{n,\alpha}$
[…]
@@
where $A_{n,\alpha}$, $B_{n,\alpha}$ are defined in (3.23)–(3.24). We estimate $B_{n,\alpha}$, which encodes the terms $\overline{\omega}\cdot\nabla u+u\cdot\nabla\overline{\omega}$. We have
@@
where $1/2=1/Q^{*}_{2}+1/p$. Second, suppose that $Q\in(1,6/5)$. (When $Q\in[6/5,2)$, then $W^{1,Q}\hookrightarrow L^{2}$, which simplifies this estimate.) Then
$\|{\rm BS}[P_{\neq 0}\omega]\cdot\nabla{\overline{\omega}}\|_{L^{6/5}}\lesssim\|{\rm BS}[P_{\neq 0}\omega]\|_{L^{Q}}^{\theta}\|{\rm BS}[P_{\neq 0}\omega]\|_{L^{2}}^{1-\theta}\|\nabla{\overline{\omega}}\|_{L^{\infty}}\,$, (3.37)
@@
$\nabla^{k+1}\omega=\mathrm{e}^{-tA_{n,\alpha}}\nabla^{k+1}\omega_{0}-\int_{0}^{t}\mathrm{e}^{-(t-s)A_{n,\alpha}}[\nabla^{k+1},A_{n,\alpha}]\omega\,ds-\int_{0}^{t}\mathrm{e}^{-(t-s)A_{n,\alpha}}\nabla^{k+1}(B_{n,\alpha}\omega)\,\mathrm{d}s\,.$ (3.41)